(no subject)

KTEV (ktev@phys.virginia.edu)
Sat, 24 Jul 1999 08:14:20 -0400 (EDT)

Shift Summary for owl shift, 24/jul/99

On Shift: Mike A., Hogan, Sami

Hours of beam delivered at all intensities: 1.1
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0.6
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:accelerator problems

========================================================================
Runs ended this shift:
none

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift:none

========================================================================
Overview of shift:
Waiting for beam until 04:00, started a run at 04:30, beam
went away again at 05:05 and returned at 06:50,
data taken in run 14010 so far is
almost entirely at low intensity and with bad targetting positions.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

0:00 to 01:30 accelerator coming up.

0:30 beam in machine but not to KTEV, clock problems
affecting magnets, also we don't get spiil
signals, making even cosmic data taking impossible.

04:00 beam returns, but is far off in horizontal (-10 wires
at NM2SEED2)

14010 04:25 begin run although beam is still mis-steered.

14010 24 05:00 beam now correctly steered, intensity about
1.8 E12.

05:07 Linac vacuum problems,

14010 110 06:50 beam returns

06:50 histo's show CsI 961 has comparator errors
reset pipeline with DBC doesn't help.
CsI 2424 has occassional large RMS (5%)
sent mail to PSh and Ed.

V0 counter 15 has low gain, mail sent to Valeri.



========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift